OpenStudy (anonymous):

Derivation of Potential energy of Magnetic field ? Can any one derive it for me

OpenStudy (anonymous):

Hi Derive potential energy of Magnetic field ?

OpenStudy (anonymous):

Potential energy of magnetic field can be derived by using the using a magnetic dipole kept in the magnetic field.  See, there is a torque which tries to align the dipole along magnetic field. The torque is \(\tau = mBsin \theta\) Work done to rotate the dipole by \(d\theta\) against the torque is \[dW = \tau d\theta = mBsin \theta d\theta\] Potential energy \( U = \int_{\theta_0 = 90^o}^{\theta} mBsin \theta d\theta = -mBcos\theta = -\vec{m}.\vec{B}\) By convention it is assumed that at \(\theta_0 = 90^o,\ U = 0\)